Calculate the heat of combustion of ethane if heat of formation is

33kJmol'

5. The heat of combustion of benzene, CoHscn and cyclohexane, CsHiżın are-

-3268 and - 3920kJmoirespectively. Calculate:

(a) the heat of hydrogenation of benzene to cyclohexane.

(5) the heat required to hydrogenate a double bond in benzene.


1
Expert's answer
2021-06-22T04:41:53-0400

Combustion of ethane is described by the equation:

C2H6(g) + 3.5 O2(g) = 2 CO2(g) + 3 H2O(l)

the heat of combustion is equal to enthalpy change in this reaction:

"\Delta"cH = - "\Delta"fH(C2H6(g)) + 2 "\Delta"fH(CO2(g)) + 3 "\Delta"fH(H2O(l))

given that formation enthaplies are "\Delta"fH(C2H6(g)) = 33 kJ/mole; "\Delta"fH(CO2(g)) = -394 kJ/mole; "\Delta"fH(H2O(l)) = -286 kJ/mole, the heat of combustion of ethane is:

Q = "\Delta"cH = -33 - 2*394 - 3*286 = -1679 kJ/mole.


1) Combustion reaction of benzene is:

C6H6(l) + 7.5 O2(g) = 6 CO2(g) + 3 H2O(l)

combustion enthaply

"\Delta"c1H = - "\Delta"fH(C6H6(g)) + 6 "\Delta"fH(CO2(g)) + 3 "\Delta"fH(H2O(l))

2) Combustion reaction of cyclohexane is:

C6H12(l) + 9 O2(g) = 6 CO2(g) + 6 H2O(g)

combustion enthaply

"\Delta"c2H = - "\Delta"fH(C6H12(l)) + 6 "\Delta"fH(CO2(g)) + 6 "\Delta"fH(H2O(l))

3) Hydrogenation of benzene is:

C6H6(l) + 3 H2(g) = C6H12(g)

(a) heat of hydrogenation of benzene to cyclohexane is equal to enthalpy in reaction 3):

"\Delta"h3H = - "\Delta"fH(C6H6(l)) + "\Delta"fH(C6H12(l)), which can be expressed from combustion enthalpies:

"\Delta"h3H = - "\Delta"fH(C6H6(l)) + "\Delta"fH(C6H12(l)) = "\Delta"c1H - "\Delta"c2H - 3 "\Delta"fH(H2O(l))

given that "\Delta"c1H = -3268 kJ/mole and "\Delta"c2H = - 3920 kJ/mole, and "\Delta"fH(H2O(g)) = -286kJ/mole, the hydrogenation enthalpy of benzene is:

"\Delta"h3H = 3268 - 3920 + 3*286= 206 kJ/mole

(b) Hydrogenation of one double bond in benzene is

C6H6(g) + H2(g) = C6H8(g)

Since the enthaply of hydogenation of benzene to cyclohexane corresponds to hydrogenation of three double bonds, the heat of hydagentaion of one bond is one-third of heat of benzene hydrogenation:

"\Delta"hH = 1/3 "\Delta"h3H = 206/3 = 68.67 kJ/mole